### v3

# Holstein mechanism in single-site model with unitary evolution

Abstract
We investigate the Holstein mechanism in a single-site model, where unitary evolution intrinsically involves both fermion and boson operators under nonadiabatic conditions. The resulting unitary dynamics and boson-frequency dependence reveal a quantum phase transition, evidenced by distinct short-time (power-law decay) and long-time (exponential decay) behaviors, which are manifested in the polaronic shift, bosonic energy, and dynamics of reduced density matrix. This observation is consistent with a non-Markovian to Markovian transition. Also, in terms of the density operator, the additional perturbation from the degenerated ground state of the total Hamiltonian would cause fluctuations that outweigh the Markovian dynamics at long times. This may lead to non-unitary evolution and highly mixed states due to entanglement with the environment.
I Introduction
Adiabaticity and non-adiabaticity fundamentally influence the physical properties of quantum systems, including their stability and phase transitions [1]. In the adiabatic limit, where the electron’s Fermi velocity significantly exceeds the sound velocity, electron-phonon scattering is elastic (even in the neutral limit), and interband transitions vanish, preventing electron energy loss. However, in the nonadiabatic regime, when a non-relativistic electron moves faster than the sound wave while still interacting with acoustic phonons, it can lose energy via Cherenkov radiation. For weak electron-phonon coupling in the stable (adiabatic) regime, the dispersion modification is linear with the impurity momentum. Conversely, for strong electron-phonon coupling, non-adiabaticity leads to an avoided crossing in the polaron band structure. Phenomena dependent on impurity motion, such as coherence or decoherence, are crucial for understanding polaron formation, as they are intrinsically linked to Fermi liquid or non-Fermi liquid behavior. For example, avoided crossings also appear in ultracold Fermi atomic systems under strong-interaction limits [6, 7, 8, 9]. Decoherence can occur in extreme cases, like a very light impurity immersed in a bath of heavy particles. Conversely, for a heavy impurity in a 1D system, decoherence can also arise at zero temperature due to the orthogonality catastrophe [2, 3, 4]. In the adiabatic case, the presence of phonon absorption and emission similarly induces decoherence, reducing coherent band motion.
In this paper, we study the single-site Holstein model. We initially choose the boson number operator with degenerated spectrum, which means the states do not distinguished by the particle number (population) but the internal degrees of freedom. When the perturbation weakly breaks the degeneracy, purity of the corresponding density matrix increases. where the internal degrees of freedom plays the main role. Also, by virtue of the high purity of the reduced density matrices, we gain the convenience of not needing to consider the deviation from the minimum uncertainty condition in the phase space.
Specifically, a time-dependent perturbation breaks the degeneracies of the original composite Hilbert spaces (electron and boson Hilbert spaces), causing the unitary evolution of the total Hamiltonian. This perturbation not only introduces time-dependence but also a Hermitian part to the boson operators. This Hermitian part of the boson operators, in turn, not only causes a nonzero commutation between the electron and boson terms but also leads the ground state of the annihilation boson operator to deviate from the classical coherent Gaussian state. The Markovian process for the system (non-adiabaticity) dominates, especially in the long-time limit. Here, the system’s relaxation time scale (characterized by exponential decay) is much slower than that of the bath electrons. Conversely, at short times, a finite non-Markovian effect with power-law decay is numerically proven, despite the absence of fluctuations as a result of non-local correlations due to the "single-site" restriction.
II Model
We consider the model described by
$H=H_{e}+H_{b}+H_{eb},$ (1)
where $H_{e}(t)=\epsilon c^{{\dagger}}(t)c(t)$ is the single electron term. Due to the single-site consideration, the bare bandwidth is irrelevant here. Note that ${\rm Tr}[(c^{{\dagger}}(t)c(t))^{2}]={\rm Tr}[c^{{\dagger}}(t)c(t)]^{2}/N$, precluding the Hubbrd-type interacion effect. We consider the total Hamiltonian in a N-by-N Hilbert space, and a time-dependent perturbation cause breaks the degeneracies of original composite Hilbert spaces. This guarantees the unitary evolution of $H(t)$, which is distinct from the Heisenberg picture. The bosonic term reads
$H_{b}(t)=\frac{p^{2}(t)}{2}+\frac{1}{2}\omega^{2}x^{2}(t)=\omega(b^{{\dagger}}(t)b(t)+\frac{{\bf I}}{2}),$ (2)
where
$b(t)=\sqrt{\frac{\omega}{2}}(x(t)+\frac{ip(t)}{\omega}),$ (3)
$b^{{\dagger}}(t)=\sqrt{\frac{\omega}{2}}(x(t)-\frac{ip(t)}{\omega}),$
$x(t)=\sqrt{\frac{1}{2\omega}}(b(t)+b^{{\dagger}}(t)),$
$p(t)=-i\sqrt{\frac{\omega}{2}}(b(t)-b^{{\dagger}}(t)),$
The boson number operator $b^{{\dagger}}(t)b(t)$ is $N$-by-$N$ Hermitian matrices with a highly degenerate spectrum and are completely degenerate initially (at t=0), reflecting the internal degrees of freedom. Also, we consider the case that the fermion number operator is always at an equilibrium steady state throughout the evolution to eliminate the potential effect of its fluctuation on the polaron shift. For electron-phonon interaction, we consider the Holstein mechanism
$H_{eb}(t)=g\omega c^{{\dagger}}(t)c(t)(b^{{\dagger}}(t)+b(t))=.g\sqrt{2\omega}\omega c^{{\dagger}}(t)c(t)x(t).$ (4)
III Results
III.1 Polaronic shift
Different to the classical harmonic oscillator where the time-dependence of boson operator can be extracted as a phase factor, which is necessary for the Lang-Firsov transformation, the boson operator here cannot. Instead, we can separate the above boson operators (non-Hermitian) into the Hermitian part and non-Hermitian part, such that
$b(t)=b_{H}(t)+b_{nH}(t),$ (5)
$b^{{\dagger}}(t)=b^{{\dagger}}_{H}(t)+b^{{\dagger}}_{nH}(t),$
$x(t)=x_{1}(t)+x_{2}(t):=\sqrt{\frac{1}{2\omega}}(b_{H}(t)+b_{H}^{{\dagger}}(t))+\sqrt{\frac{1}{2\omega}}(b_{nH}(t)+b_{nH}^{{\dagger}}(t)),$
$p(t)=p_{1}(t)+p_{2}(t):=-i\sqrt{\frac{\omega}{2}}(b_{H}(t)-b_{H}^{{\dagger}}(t))-i\sqrt{\frac{\omega}{2}}(b_{nH}(t)-b_{nH}^{{\dagger}}(t)),$
where $[b_{H}(t),b^{{\dagger}}_{H}(t)]=0$ and $[b_{nH}(t),b^{{\dagger}}_{nH}(t)]=1$. Here the boson operators are still non-Hermitian and the position/momentum operators are still Hermitian. Note that ${\rm Tr}[x_{1}(t)^{2}]={\rm Tr}[x_{1}(t)]^{2}/N$ and ${\rm Tr}[x_{2}(t)^{2}]={\rm Tr}[x_{2}(t)]^{2}/N$, i.e., $x_{1}(t)$ and $x_{2}(t)$ are diagonally dominant matrices. We consider the case where the phase-space distribution remains, and as a product of the Gaussian Wigner distributions $x^{2}$ and $p^{2}$ which are inversely proportional, i.e., $\langle x^{2}\rangle=(4\langle p^{2}\rangle)^{-1}$, for the case of squeezed state at zero-temperature. This minimum uncertainty is guaranteed by the fact that both the total density matrix and the reduced one are all pure state during the evolution. Next we introduce the anti-Hermitian operator (such that $S^{{\dagger}}=-S$ and $e^{S}$ is unitary)
$S=c^{{\dagger}}(t)c(t)(b^{{\dagger}}(t)-b(t))=\frac{\sqrt{2}}{i\sqrt{\omega}}c^{{\dagger}}(t)c(t)p(t)$ (6)
and apply the Lang-Firsov transformation $\overline{O}=e^{S}Oe^{-S}$ to $H$, we obtain
$\overline{H}=\epsilon c^{{\dagger}}(t)c(t)+\omega b^{{\dagger}}(t)b(t)+\frac{{\bf I}}{2}+c^{{\dagger}}(t)c(t)\Sigma(t),$ (7)
where only the fermion number operator is invariant under the transformation. $\Sigma(t):=\Sigma_{eb}(t)+\Sigma_{b}(t)=g\omega e^{S}(b^{{\dagger}}(t)+b(t))e^{-S}+(e^{S}H_{b}e^{-S}-H_{b})H_{b}^{-1}$ is the polaronic self-energy. This definition is consistent with Ref.[5], i.e., the transformed fermion-boson term and the differece between the fermion terms after and before transformation, which measures the shift of bosonic oscillator due to the presence of electron. Note that all the expectations in this article are in the single-electron basis state $c^{{\dagger}}(t)|0\rangle$. All the numerical simulations perfomed in this article base on a 2-by-2 electron Hilbert space and 4-by-4 boson Hilbert space, such that there is 8-dimensional combined basis ($N=8$), and we adopt the convenience of notation $c^{{\dagger}}\equiv c^{{\dagger}}_{2\times 2}\otimes{\bf I}_{4\times 4}$.
Due to the intrinsic unitarility we have ${\rm Tr}H_{b}(t=0)=N\omega$ and ${\rm Tr}H_{b}(t=\infty)=\omega$, where there is a smooth exponential decay with time. For $e^{S}x(t)e^{-S}$, diagonalization through the transformation of $e^{S}$ is only successful on $x_{2}(t)$,
$\frac{1}{g\omega N}{\rm Tr}[\Sigma_{eb}(0)]=\frac{1}{N}{\rm Tr}e^{S}(b^{{\dagger}}(t)+b(t))e^{-S}=\frac{1}{N}{\rm Tr}e^{S}(b^{{\dagger}}_{H}(t)+b_{H}(t))e^{-S}+\frac{1}{N}{\rm Tr}e^{S}(b_{nH}^{{\dagger}}(t)+b_{nH}(t))e^{-S}$ (8)
$\sim\left\{\begin{array}[]{ll}2-{\rm ln}\omega,\ (\omega<1)\\ 2+(e^{-(\omega-1)}-1)\ (\omega>1)\end{array},\right.$
where the factor 2 originates from ${\rm Tr}[b_{H}(t=\infty)+b_{H}^{{\dagger}}(t=\infty)]=2$ (${\rm Tr}[b_{H}(t=0)+b_{H}^{{\dagger}}(t=0)]=2N$). The time and frequency dependence are important to seeking the minimal of polaronic shift. For $\omega\leq 1$ and at finite time,
$\frac{1}{g\omega N}{\rm Tr}[\Sigma_{eb}(t)]=\frac{1}{N}{\rm Tr}e^{S}(b^{{\dagger}}(t)+b(t))e^{-S}=\frac{1}{N}{\rm Tr}e^{S}(b^{{\dagger}}_{H}(t)+b_{H}(t))e^{-S}+\frac{1}{N}{\rm Tr}e^{S}(b_{nH}^{{\dagger}}(t)+b_{nH}(t))e^{-S}$ (9)
$\sim\left\{\begin{array}[]{ll}(2-{\rm ln}\omega)\frac{\omega}{t},\ (small\ t)\\ (2-{\rm ln}\omega)e^{-t/\omega}\ (large\ t)\end{array}\right.+\left\{\begin{array}[]{ll}(2-{\rm ln}\omega)\frac{\omega}{t^{1/\omega}},\ (small\ t)\\ -\frac{(2-{\rm ln}\omega)e^{-t/\omega}+1}{\omega}\ (large\ t)\end{array},\right.$
where the early stage and latter stage follow the power-law decay and exponential decay, respectively. This is shown in Fig.2, in terms of the boson dispalcement. The expectation of $H_{b}$ in total density
$\frac{1}{\omega}\langle H_{b}\rangle$$\sim\left\{\begin{array}[]{ll}\frac{1}{t},\ (small\ t)\\ e^{-t}\ (large\ t),\end{array}\right.$ (10)
as shown in Fig.1(a). Thus, the expectation exhibits a transition from power-law decay to exponential decay with time evolution. This signifies that the corresponding system density’s dynamics (Lindbladian) is gapless in the beginning and becomes gapped at long times. Note that during the whole process, the spectrum of the Lindbladian should be purely real (without the component of oscillation frequencies) due to the absence of non-local correlations and coherence (electron hopping) for $\omega\leq 1$.
Figure 1: (a) Expectation of boson operator on the single electron basis state. at long-time. (b)-(d) Expectations of position and momentum operators which follow the Gaussian Wigner distributions. (b) shows the case $\omega>1$ where the imaginary part of squared position and momentum operators emerges (this is also the regime where there is nonzero real part of expectation of momentum operator).
Figure 2: Expectations of $x_{1}(t)$ and $x_{2}(t)$. The orange and green lines fit the early and latter stages using power-law and exponential decay, respectively.
Figure 3: Expectations of $\Delta x^{2}$ as a function of time ((a)-(b)) and frequency ((c)). Above the critical frequency $\omega_{c}=1$, the real part star to decline and nonzero imaginary part emerges. (d)The same with (a)-(b) but replacing the $S$ by $S^{\prime}$.
III.2 Effect of transformation to the boson operators
Since both the fermion operator and boson operator share the same unitary dynamics, they are not mutually commute unless remove the Hermitian part of the boson operators. The difference of position operator before and after transformation, $\sqrt{2\omega}\Delta x(t):=e^{S}(b^{{\dagger}}(t)+b(t))e^{-S}-(b^{{\dagger}}(t)+b(t))$ which is a diagonalizable defective matrix, is not soly determined by the fermion operator due to the contribution from $[b^{{\dagger}}_{nH}(t)-b_{nH}(t),b^{{\dagger}}_{H}(t)+b_{H}(t)]$. The lowered rank signifies the existence of states insensitive to the perturbations (reminiscent of the dark states). Further, since $e^{S}(b^{{\dagger}}(t)+b(t))^{2}e^{-S}=((b^{{\dagger}}(t)+b(t))+\sqrt{2\omega}\Delta x(t))^{2}$, we have
${\rm Tr}[2\omega\Delta x(t)^{2}]+2{\rm Tr}[(b^{{\dagger}}(t)+b(t))\sqrt{2\omega}\Delta x(t)]=0,$ (11)
where for $\omega\leq 1$, we obtain
${\rm Tr}[2\omega\Delta x(t)^{2}]\sim\left\{\begin{array}[]{ll}0,\ (small\ t)\\ \omega(1-e^{-t})\ (large\ t)\end{array}\right.$ (12)
which is linear in $\omega$ at long time. The exact result of ${\rm Tr}[2\omega\Delta x(t)^{2}]$ is shown in Fig.3. We also obtain its approximated result in appendix. Fluctuations signifying non-local correlations can be seen in Fig.3(d), where we apply another transformation using $S^{\prime}=c^{{\dagger}}(0)c(t)(b^{{\dagger}}(t)-b(t))$, which is nomore a anti-Hermitian operator as long as $t\neq 0$. Nonzero value of ${\rm Tr}[2\omega\Delta x(t)^{2}]$ also implies the squeezing effect from the perturbation, similar to the result of canonical transformation for squeezing with squeezing operator $e^{r(b^{2}-(b^{{\dagger}})^{2})}$.
IV Dynamics of reduced density matrix
We next focus on the dynamics of reduced density matrix $\rho_{e}={\rm Tr}_{b}\rho_{tot}$ (partial trace over the bosonic bases) where $\rho_{tot}=|\psi_{0}\rangle\langle\psi_{0}|$ with $|\psi_{0}\rangle$ the eigenstate of $H$ corresponding to lowest eigenvalue, i.e., $\rho_{tot}=\sum_{ij}c_{i}c_{j}^{*}|\Psi_{i}\rangle\langle\Psi_{j}|$ where $c_{i}$ is the $i$-th component of $|\psi_{0}\rangle$, and $|\Psi_{i}\rangle$ is the $i$-th state of the combined basis. Note that the rank of $H$ is lowered with increasing time, in this case we use $\rho_{tot}=\frac{1}{d}\sum_{j=1}^{d}|\psi_{0;j}\rangle\langle\psi_{0;j}|$ with $d$ the ground state degeneracy, in which case it requires further orthonormalized to keeping ${\rm Tr}\rho_{e}=1$. While for $\rho_{tot}$ in mixed state which corresponds to the case of finite temperature, we have $\rho_{tot}=\sum_{j}e^{-\beta E_{j}}|\psi_{j}\rangle\langle\psi_{j}|/\sum_{j}e^{-\beta E_{j}}$. But we stick to the zero temperature case in this paper, thus we do not adopt this mixed state. The resulting reduced electron density $\rho_{e}$ is a 2-by-2 Hermitian matrix, with the two diagonal elements describe the populations of two states in electron Hilbert space, and the two off-diagonal elements describe the coherence between the two states. The result shows that, at short-time (far before the emergence of degenerated ground state of $H$), $\rho_{e;11}\sim{\rm ln}t(=1-\rho_{e;22})$, and reasonably, $\frac{d\rho_{e;11}}{dt}\sim\frac{1}{t}(=-\frac{d\rho_{e;22}}{dt})$, as shown in Fig.4. The logarithmic increase in population $\rho_{e;11}$ reflects a extremely slow dynamics, and the system is not reaching a steady state. The power-law decay in $\frac{d\rho_{e;11}}{dt}$ again signifying the Non-Markovian Memory effect in early stage during the evolution. While exponential decay emerges at longer time (but before the emergence of degenerated ground state of $H$), $\rho_{e;11}\sim-e^{-t}$ and $\frac{d\rho_{e;11}}{dt}\sim e^{-t}$.
As shown in Fig.4, $\rho_{e;11}$ is quite large and increase with time, thus $\rho_{e}$ and $\rho_{b}$ are close to purity state and the purity increase with time. This is consistent with the increased purity of each subsystem as shown in Fig.5(d) and the continously decrease of internal entanglement (negativity) of $\rho_{tot}$ as shown in Fig.6(b). Thus the nonMarkovian-to-Markovian transition can be well observed at least before the critical time $t_{c}$ of the emergent degenerated ground state of $H$, during which the evolution is unitary and the system is nearly closed such that as the total density getting thermalized by the environment with lowered internal entanglement, the subsystems have increased purity until the entanglement with the environment sets in (the begining of nonunitary evolution). In other word, in early stage, the weak nonMarkovianity leads to power-law monotonous decay; while in late stage, the degenerate ground state of $H$ cause additional strong nonMarkovianity with significant coherent exchange of information between system and bath, which outweights the Markovian thermal process despite the decoherence dominates at the long-time and implying the enhanced entanglement. Note that the Markovian dynamics in the late stage can still be observed in term of the observables regardless of the density operators. Thus, the perturbation, which is initially well-defined and coherent, in the late stage becomes random and is composed of many unobserved degrees of freedom due to the degeneracy of the system’s ground state.
The dissipative phase transition is only observable in the early stages of evolution. During this period, the dissipation-induced gain and loss can be ignored, and the total Hamiltonian maintains unitary evolution. By setting a finite threshold value for the ground state, Fig.5 presents the scenario where the total density matrix is considered to be in a mixed state. In this case, once time exceeds a critical value, the total system’s density matrix becomes mixed. This implies that the system is indeed an open system and its evolution becomes non-unitary at long times. Here, strong dissipation drives the system toward a state akin to a highly fluctuating thermal state in a closed system (see Fig.5(b)). This observation is consistent with the opening of the dissipative gap in the later stage of evolution. At this point, quantum fluctuations (under the zero-temperature condition) make continuous symmetry breaking and the formation of an ordered phase less likely.
Figure 4: First diagonal element of the reduced density matrix $\rho_{e}$ and its time derivative. The vertical dashed line indicate the time where the degenerated ground states of $H$ emege.
Figure 5: (a) The four diagonal elements (populations) of $\rho_{b}$ for unitary evolution and pure total density, obtained by setting an infinitesimally small threshold value for judging the degenerate ground state. (b)-(c) Populations of $\rho_{b}$ and $\rho_{e}$ when a finite threshold value is used. The non-unitary evolution and mixed total density appear at a critical time $t_{c}\approx 800$, as indicated by the vertical cyan dashed line. (d) Purities of $\rho_{tot}$ (blue) and $\rho_{e}$ (or $\rho_{b}$; red) for finite threshold value (left axis) and infinitesimally threshold value (right axis). The green dashed line indicates when the population starts to deviate from its logarithmic increase, i.e., the emergence of Markovian process. The left red arrow indicates the lowered purity due to the entanglement between subsystems, while the right red arrow indicates the lowered purity due to the entanglement between the subsystem and the bath. Obviously, the effect of the latter dominates over the fomer. That is why the nonMarkovian-type fluctuation overweight the Markovian dynamics in the late stage.
V Conclusion
[…]
The unitary dynamics preserves the phase-space distribution during the Markovian process wth slow relaxation, and the electron-phonon interaction dominates over the influences from bath electrons. In this limit, the Gaussian wave function fail to describe bosons which can no longer be treated as classical field, and phononic time scale is much faster than the electronic time scale. The mean-field theory also fails to be applied in this case. The unitary dynamics as well as boson-frequency-dependence provides the evidence of quantum phase transition from the distinct behaviors at short-time and long-time stages, which exhibit power law and exponential law decay, respectively, as can be seen from the polaronic shift, boson energy, and the dynamics of reduced density $\rho_{e}$. Moreover, the effective mass renormalization $m^{*}/m(=e^{\Sigma(t)/\omega})$ should follow the same rule, despite not being shown here. The gradual opening of the dissipative gap can be further verified by estimating the Lindbladian’s spectrum. This estimation, achievable through Lanczos iteration or the quantum trajectory method, as well as the case of boson squeezing ($(b^{{\dagger}})^{2}+b^{2}$) and the deviation away from minimum uncertainty condition (entanglement increase the classical uncertainty), will be the scope of our next work.